Kingston Company uses the dollar-value LIFO method of computing inventory. An external price index is used to convert ending inventory to base year. The company began operations on January 1, 2021, with an inventory of $201,000. Year-end inventories at year-end costs and cost indexes for its one inventory pool were as follows:

Year Ended Ending Inventory Cost Index
December 31 at Year-End Costs (Relative to Base Year)
2021 $ 291,600 1.08
2022 376,420 1.18
2023 355,350 1.15
2024 349,650 1.11

Calculate inventory amounts at the end of each year

Inventory Layers Converted to Base Year Cost Inventory Layers Converted to Cost Inventory DVL Cost
Date Inventory at Year-End Cost Year-End Cost Index Inventory Layers at Base Year Cost Inventory Layers at Base Year Cost Year-End Cost Index = Inventory Layers Converted to Cost
01/01/2021 = Base =
12/31/2021 = Base =
2021 =
12/31/2022 = Base =
2021 =
2022 =
12/31/2023 = Base =
2021 =
2022 =
2023 =
12/31/2024 = Base =
2021 =
2022 =
2023 =
2024 =
Answer #1

Calculate inventory amounts at the end of each year

Inventory Layers Converted to Base Year Cost Inventory Layers Converted to Cost Inventory DVL Cost
Date Inventory at Year-End Cost Year-End Cost Index Inventory Layers at Base Year Cost Inventory Layers at Base Year Cost Year-End Cost Index = Inventory Layers Converted to Cost
01/01/2021 201000 / 1 = 201000 Base 201000 * 1 = 201000
12/31/2021 291600 / 1.08 = 270000 Base 201000 * 1 = 201000
2021 69000 * 1.08 = 74520 275520
12/31/2022 376420 / 1.18 = 319000 Base 201000 * 1 = 201000
2021 69000 * 1.08 = 74520
2022 49000 * 1.18 = 57820 333340
12/31/2023 355350 / 1.15 = 309000 Base 201000 * 1 = 201000
2021 69000 * 1.08 = 74520
2022 39000 * 1.18 = 46020 321540
2023 =
12/31/2024 349650 / 1.11 = 315000 Base 201000 * 1 = 201000
2021 69000 * 1.08 = 74520
2022 39000 * 1.18 = 46020
2023 0 * 0 = 0
2024 6000 * 1.11 = 6660 328200